OpenStudy (anonymous):

Simplify the expression √96 over 8

OpenStudy (lukecrayonz):

sqrt(3/2)

OpenStudy (anonymous):

that is not in my choices

OpenStudy (lukecrayonz):

What are your choices?

OpenStudy (anonymous):

sqrt23/4

OpenStudy (anonymous):

the choices are: A. \[4\sqrt{3} \over \sqrt{2}\] B. \[\sqrt{2} \over 4\] C. \[2\sqrt{3}\] D. \[1 \over 2\sqrt{3}\]

OpenStudy (mertsj):